A projectile is fired horizontally from a 40 m high roof of a building and lands 85 m from the base of the building. What is the initial velocity of the projectile?
Answer #1

firstly the projectile is fired horizontally from a height which means that height is tha maximum height of that projectile so,

H = 0.5*g*t²

where H is the height of building g is the acceleration due to gravity and t is time taken.

now putting all the given values,

40 = 0.5 * 9.8 * t²

solving for t,

t = √ 8.16

t = 2.86 s

The key to these type of questions is finding t first like this because it is the same for the vertical and horizontal components.

Since acceleration a is zero in the horizontal component this becomes, S = u * t

where S is the distance traveled in the horizontal direction and u is the initial velocity of the projectile.

85 = u * 2.86

u = 29.7 m/s

so, the initial velocity of the projectile is 29.7 m/s.

→ if you find this answer helpful please rate positive thank you so much.